which meathod would you use to solve the following system? please explain.
3x+2y=9
-2x+3y=5
thanks for all the help. it is well appricated :)

Answer by stanbon(75887)

which meathod would you use to solve the following system? please explain.
3x+2y=9
-2x+3y=5
------
Modify for elimination to solve for "y":
-------
6x + 4y = 18
-6x+ 9y = 15
--------
Add and solve for "y":
13y = 33
y = 33/13
-----
Modify the original set to solve for "x":
9x + 6y = 27
-4x + 6y = 10
----------------
Subtrac to solve for "x":
13x = 14
x = 14/13
------------------
Cheers,
Stan H.